What is the perimeter of the smallest box that this record can fit into with a circumference of 8 inches? URGENT!!!
Answer by Theo(3504)

you want to put a circle inside a box.

you are circumscribing the circle with the box.

you are inscribing the box with the circle


the circumference of your circle is 8 inches.

you need to find the diameter.

the circumference of a circle is equal to 2 * pi * r.

your equation to solve is 2 * pi * r = 8

divide both sides of the equation by 2 * pi to get r = 8 / (2 * pi)

pi = 3.141592654...

this is a constant term that never changes.

using that, we find that r = 1.273239545

that's the radius of your circle.

you want to get the diameter of your circle, so you multiply the radius by 2 to get:

diameter of the circle equals 2.546479089 inches.

that would be the width of your box.

the box would be square.

the perimeter of your box would be equal to 4 * 2.546479089 = 10.18591636 inches.